ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 53C20, 53C30, 53C35, 53B20 Nikčević, Ž. Stana, Mathematical Institute, SANU, Belgrade, Yugoslavia; Kowalski, Oldřich, Charles University, Praha, Czech Republic Curvature Homogeneous Spaces and Related Topics We study contact of higher order of Riemannian manifolds (or metrics). The main results concern curvature homogeneous Riemannian manifolds of arbitrary order and envelopes of Riemannian metrics of the 1st and 2nd order.